Managerial AccountingImperial Jewelers manufactures and sells a gold bracelet for $406.00. The companys accounting system says that the unit prod
Answer #1

Special order revenue = 22*366 = 8052

Cost of making = (146+10+87+11)*22 + 455 = 6043

Financial advantage = 2009

(8052 - 6043)

2. Yes